Before hanging new William Morris wallpaper in her bedroom, Brenda sanded the walls lightly to smooth out some

irregularities on the surface. The sanding block weighs 1.90 N and Brenda pushes on it with a force of 3.00 N at an angle

of 30.0° with respect to the vertical, and angled toward the wall.What is the coefficient of kinetic friction between the wall and the block?


1
Expert's answer
2022-03-14T09:45:57-0400

"\vec P+\vec F+\vec F_{fr}=\vec 0,"

"P-F\sin30°=\mu P\cos 30°,"

"\mu=\frac{P-F\sin30°}{P\cos 30°}=0.24."